# StructFlowBench: A Structured Flow Benchmark for Multi-turn Instruction Following

### v1

@@ Abstract
Multi-turn instruction following capability constitutes a core competency of large language models (LLMs) in real-world applications. Existing evaluation benchmarks predominantly focus on fine-grained constraint satisfaction and domain-specific capability assessment, yet overlook the crucial structural dependency between dialogue turns that distinguishes multi-turn from single-turn interactions. This structural dependency not only reflects user intent but also establishes a second dimension for instruction following evaluation beyond constraint satisfaction. To address this gap, we propose StructFlowBench, a multi-turn instruction following benchmark with structural flow modeling. The benchmark innovatively defines a structural flow framework comprising six fundamental inter-turn relationships, which not only introduces novel structural constraints for model evaluation but also serves as generation parameters for creating customized dialogue flows tailored to specific scenarios. Adopting established LLM-based automatic evaluation methodologies, we conduct systematic evaluations of 13 leading open-source and closed-source LLMs. Experimental results reveal significant deficiencies in current models’ comprehension of multi-turn dialogue structures. The code is available at https://github.com/MLGroupJLU/StructFlowBench.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) in multi-turn dialogue systems has elevated instruction-following capabilities to a pivotal research frontier in human-AI interaction Chang et al. (2024). Current evaluation methodologies bifurcate into two streams: multi-turn dialogue evaluations focusing on capability evaluation Zheng et al. (2023b); Bai et al. (2024); Kwan et al. (2024) and instruction-following analyses emphasizing fine-grained constraint compliance Jiang et al. (2024); He et al. (2024a); Zhang et al. (2024). More recent research has started to model the composition of intra-turn constraints Wen et al. (2024).
However, current evaluation methodologies treat multi-turn dialogues as simple concatenations of single-turn interactions, neglecting users’ planning and intentionality in extended conversations. This leads to three critical limitations: (1) Failure to model complex scenarios: Multi-turn dialogue data constructed with simplistic linear thinking cannot accurately capture key characteristics of real-world complex conversations, such as logical coherence, user goal clarity, and natural transitions. (2) Methodological bias: Single-turn evaluation strategies fragment inter-turn structural connections, overlooking multi-turn structural constraints. (3) Analytical deficiency: Existing approaches overemphasize intra-turn-level constraint compliance while lacking a systematic framework to characterize dialogue structural flow.
@@
To bridge these gaps, we introduce StructFlowBench, a novel instruction-following benchmark integrating a multi-turn structural flow framework. It consists of two key components: 1) Dual-constraint evaluation system, which combines 8 intra-turn instruction constraints with 5 newly proposed structural constraints, enabling a more comprehensive assessment of multi-turn dialogue instruction following capabilities of LLMs. These structural constraints account for inter-turn dependencies, ensuring that models are evaluated not only on their ability to satisfy individual constraints but also on their capacity to maintain logical coherence across multiple turns. 2) Six-category structural flow taxonomy, encompassing six fundamental inter-turn relationships: Follow-up, Refinement, Recall, Summary, Expansion, Unrelatedness. The illustration of the structural flow taxonomy and an example of the structural flow are presented in Figure 1. This taxonomy serves a tripartite function: (a) Diagnostic evaluation: It enables a structured analysis of cross-turn structural rationality, helping to identify inconsistencies in dialogue flow and ensuring that model responses align with the expected discourse structure. (b) Intent inference: By analyzing structural patterns, this taxonomy facilitates the extraction of implicit user intent, offering a deeper understanding of how instructions evolve over multiple turns. (c) Controlled generation: The taxonomy provides configurable structural parameters that guide task-specific dialogue simulation, allowing for the tailored generation of multi-turn conversations with predefined structural patterns. This not only enhances dataset diversity but also supports the development of more robust instruction-following models adaptable to varied real-world applications.
@@
  • Structural Flow Taxonomy: We propose a six-category structured taxonomy for multi-turn instruction-following evaluation, offering an interpretable framework for analyzing dialogue structural flow.
  • StructFlowBench: We introduce StructFlowBench, a structurally annotated multi-turn benchmark that leverages a structure-driven generation paradigm to enhance the simulation of complex dialogue scenarios.
@@ 2.1 Benchmarks for Multi-Turn Dialogues
The evolution of dialogue evaluation paradigms has progressed from single-turn assessments to sophisticated multi-turn interaction analysis Wang et al. (2023); Sun et al. (2024); Duan et al. (2024). Among these, MT-Bench Zheng et al. (2023b) pioneered this transition by providing methodologies specifically designed to assess a model’s ability to handle multi-turn interactions. Building upon this, MT-Bench-101 Bai et al. (2024) introduces a more granular evaluation framework to assess fine-grained capabilities. Multi-IF He et al. (2024b) expands single-turn dialogues into multi-turn interactions by following simple, predefined linear paths. However, most existing work on multi-turn dialogue evaluation does not prioritize the assessment of instruction following and overlooks the influence of structural information on the evaluation of multi-turn dialogues. The four modes of user-assistant interactions proposed by MT-Eval Kwan et al. (2024) cover certain structural information within multi-turn dialogues. However, MT-Eval does not establish a systematic structural framework and lacks integration of various structural aspects for a comprehensive evaluation.
@@ 2.2 Benchmarks for Instruction Following
Recent instruction following evaluation predominantly employs constraint-based frameworks Jiang et al. (2024); Zhang et al. (2024); He et al. (2024a); Zhou et al. (2023). InfoBench Qin et al. (2024) introduces the Decomposed Requirements Following Ratio (DRFR) metric, which provides a more granular scoring system by breaking down the evaluation of complex instructions into assessments of their individual simple constraints. Furthermore, ComplexBench Wen et al. (2024) explores instruction-following capabilities in single-turn complex dialogues through empirical studies of constraint composition. However, prior work on instruction-following evaluation has primarily focused on single-turn interactions, which do not align with the more common multi-turn dialogue scenarios observed in real-world user interactions. While some studies have attempted to split complex single-turn instructions into multi-turn dialogues, these approaches do not fully capture the intentionality and goal-oriented nature of users in real-world contexts.
@@ 3 StructFlowBench
In this section, we first introduce the structural flow framework and the constraint categories of our benchmark. Next, we detail the data construction pipeline and present an overview of the statistics for StructFlowBench. Finally, we outline the evaluation protocol.
@@ 3.1 Structural Flow Taxonomy
By analyzing existing LLM and real human multi-turn dialogue datasets (such as WILDCHAT Zhao et al. (2024) and LMSYS-Chat-1M dataset Zheng et al. (2023a)), we identified and categorized six structural patterns of multi-turn dialogues to enhance the understanding and analysis of conversational structural flow.
Follow-up: An adjacent-turn structure where the user’s next prompt builds on the content of the previous turn, incorporating details from either the user’s previous prompt or the AI’s previous response. This is the most common structure in multi-turn dialogues, typically reflecting the user’s intent to explore the topic more deeply.
Refinement: An adjacent-turn structure in which the user modifies or clarifies their immediate previous prompt to improve the AI’s response. This structure usually signals the user’s dissatisfaction with the prior response, prompting them to refine the prompt while clarifying and emphasizing their concerns to obtain a more satisfactory response.
Recall: A long-range structure in which the user refers back to content from two or more turns ago to provide context for the current prompt (long-range follow-up) or referencing prior content for clarification (long-range refinement).
@@
After defining the six basic dialogue structures, we can use the Structural Flow Taxonomy to analyze multi-turn dialogue data and construct the corresponding structural flows.
@@ 3.2 Constraint Categories
We categorize our constraints into intra-turn constraints and multi-turn structural constraints. Details related to constraints can be found in Appendix B.
For intra-turn constraints, we synthesize and refine constraint classification systems from several works in this field (e.g., IF-Eval Zhou et al. (2023), CFBench Zhang et al. (2024), FollowBench Jiang et al. (2024)). Based on this synthesis, we categorize constraints into eight types: Inverse Constraint, Style Constraint, Situation Constraint, Keyword/Element Constraint, Basic Format Constraint, Quantity Format Constraint, Template Format Constraint, and Content Constraint.
For multi-turn structural constraints, we define five types of structural design constraints, excluding the “unrelatedness” structure. These constraints are specifically designed to maintain logical coherence and continuity across multiple turns in a dialogue. They ensure that the structural relationships between turns are consistent and contextually relevant, enabling a smooth flow of conversation. The five types of constraints are aimed at handling key aspects such as follow-ups, refinements, recalls, expansions, and summaries, ensuring that each turn in the dialogue properly connects to the previous ones while adhering to the intended conversational structure.
@@ 3.3 Data Construction Pipeline
The construction pipeline of StructFlowBench, as shown in Figure 2, comprises three main components: parameter setting, two-step dialogue generation, and constraint extraction and addition. All prompt templates used in the data construction process are included in Appendix D, and a sample data instance is provided in Appendix E.
@@
Before dialogue generation, we select parameters such as topic, task, user characteristics, and structural flow template, ensuring comprehensive coverage of the evaluation scope for multi-turn dialogue generation. For task types, we refer to the taxonomy of ComplexBench Wen et al. (2024), adapting it to our evaluation framework and selecting eight task types. For topics, we draw from the MT-Bench-101 Bai et al. (2024) framework, making necessary adjustments to suit our context, and ultimately select 22 topics. For user characteristics, we consider the significant differences in questioning styles and language between experts and non-experts. For the structural flow template, we designed multiple templates based on insights from real data and specific scenarios.
Two-Step Dialogue Generation
We employ a two-step process to generate a dialogue for a parameter setting. The first step uses the structural flow template to generate an intermediate dialogue plan (i.e., summarized prompts) via GPT-4o. Locally deployed mini-models perform initial screening and manual inspection of error data to ensure the dialogue plan aligns with the structural flow. In the second step, each intermediate dialogue plan is used to generate a complete dialogue, including user prompts and LLM responses via GPT-4o. This approach ensures high-quality generation of both dialogue content and structure while minimizing manual effort.
Constraint Extraction and Addition
For the complete multi-turn dialogue data, we extract intra-turn constraints using the GPT-4o, followed by manual validation to ensure accuracy. Based on the structural flow information, we then assign the corresponding multi-turn structural constraints to each dialogue turn.
3.4 Benchmark Dataset Statistics
@@
Table 1: Comparisons between StructFlowbench and other related benchmark datasets. $\triangle$ represents partially satisfied.
@@ 3.5 Evaluation
Deepseek-v3 0.99 0.8 0.92 1.0 1.0 0.98 0.93 0.96 0.98
Gemini-1.5-Pro 0.97 0.78 0.91 1.0 0.94 0.97 0.91 0.95 0.97
GPT-4o 0.98 0.78 0.88 0.97 0.91 0.97 0.9 0.95 0.97
Claude-3.5-Sonnet 0.98 0.8 0.88 1.0 0.91 0.95 0.88 0.94 0.96
GLM-4-9B-Chat 0.95 0.75 0.84 0.97 0.94 0.95 0.86 0.93 0.95
Qwen2.5-14B-Instruct 0.97 0.73 0.87 0.97 0.97 0.94 0.84 0.92 0.94
Qwen2.5-7B-Instruct 0.95 0.76 0.9 0.94 0.97 0.94 0.84 0.92 0.94
Deepseek-R1-Distill-Qwen-7B 0.91 0.62 0.85 0.86 0.78 0.81 0.69 0.8 0.82
@@
Llama-3.1-Instruct-8B 0.96 0.71 0.84 0.79 0.94 0.85 0.68 0.83 0.86
Phi-3.5-mini-instruct 0.94 0.68 0.87 0.94 0.94 0.88 0.73 0.87 0.88
@@
Mistral-7B-Instruct-v0.3 0.97 0.59 0.87 0.71 0.97 0.77 0.56 0.76 0.78
Table 2: StructFlowBench rated by GPT-4o. The left side displays the performance of various models on the five basic structural constraints, while the right side presents their performance on the four key metrics.
@@ 4.1 Experimental Setup
We evaluate 13 popular LLMs on StructFlowBench, including 3 closed-source models (GPT-4o Hurst et al. (2024), Claude-3.5-Sonnet Anthropic (2024) and Gemini-1.5-Pro Team et al. (2024) ) and 10 open-source models: Llama-3.1-Instruct-8B Dubey et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct Yang et al. (2024), Yi-6B-Chat Young et al. (2024), Phi-3.5-mini-instruct Abdin et al. (2024), GLM-4-9B-Chat GLM et al. (2024), Deepseek-R1-Distill-Llama-8B, Deepseek-R1-Distill-Qwen-7B Guo et al. (2025) and DeepSeek-v3 Liu et al. (2024). More details on these evaluated models can be found in Appendix F.

### v2

@@ Abstract
Multi-turn instruction following capability constitutes a core competency of large language models (LLMs) in real-world applications. Existing evaluation benchmarks predominantly focus on fine-grained constraint satisfaction and domain-specific capability assessment, yet overlook the crucial structural dependencies between dialogue turns that distinguish multi-turn from single-turn interactions. These structural dependencies not only reflect user intent but also establish an essential second dimension for the instruction following evaluation beyond constraint satisfaction. To address this gap, we propose StructFlowBench, a multi-turn instruction following benchmark with structural flow modeling. The benchmark defines an innovative structural flow framework with six fundamental inter-turn relationships. These relationships introduce novel structural constraints for model evaluation and also serve as generation parameters for creating customized dialogue flows tailored to specific scenarios. Adopting established LLM-based automatic evaluation methodologies, we conduct systematic evaluations of 13 leading open-source and closed-source LLMs. Experimental results reveal significant deficiencies in current models’ comprehension of multi-turn dialogue structures. The code is available at https://github.com/MLGroupJLU/StructFlowBench.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) in multi-turn dialogue systems has elevated instruction-following capabilities to a pivotal research frontier in human-AI interaction Chang et al. (2024). Current evaluation methodologies bifurcate into two streams: multi-turn dialogue evaluations focusing on general capabilities Zheng et al. (2023b); Bai et al. (2024); Kwan et al. (2024) and instruction-following analyses emphasizing fine-grained constraint compliance Jiang et al. (2024); He et al. (2024a); Zhang et al. (2024). More recent research has started to model the composition of intra-turn constraints Wen et al. (2024).
@@
Figure 2: An example of the Structural Flow demonstrates how inter-turn relationships are represented using the proposed taxonomy, offering insight into the structure of real dialogues.
However, current evaluation methodologies treat multi-turn dialogues merely as simple concatenations of single-turn interactions, neglecting users’ planning and intentionality in extended conversations. This leads to three critical limitations: (1) Failure to model complex scenarios: Multi-turn dialogue data constructed with simplistic linear thinking cannot accurately capture key characteristics of real-world complex conversations, such as logical coherence, user goal clarity, and natural transitions. (2) Methodological bias: Single-turn evaluation strategies fragment inter-turn structural connections, overlooking multi-turn structural constraints. (3) Analytical deficiency: Existing approaches overemphasize intra-turn constraint compliance while lacking a systematic framework to characterize dialogue structural flow.
To bridge these gaps, we introduce StructFlowBench, a novel instruction-following benchmark integrating a multi-turn structural flow framework. It consists of two key components: 1) Dual-constraint evaluation system, comprising two layers—intra-turn constraint evaluation and inter-turn constraint evaluation—offers a more comprehensive assessment of LLMs’ multi-turn dialogue instruction-following capabilities. Building on existing intra-turn constraint assessments, we integrate five newly proposed structural constraints with eight intra-turn instruction constraints to capture the relationships between dialogue turns. These structural constraints account for inter-turn dependencies, ensuring that models are evaluated not only on their ability to satisfy individual constraints but also on their capacity to maintain logical coherence across multiple turns. 2) Six-category structural flow taxonomy, encompassing six fundamental inter-turn relationships: Follow-up, Refinement, Recall, Summary, Expansion, Unrelatedness. Figure 1 illustrates the structural flow taxonomy, which is defined in detail in Section 3.1. This taxonomy serves a tripartite function: (a) Structural Diagnosis: It enables a structured analysis of cross-turn structural rationality, helping to identify inconsistencies in dialogue flow and ensuring that model responses align with the expected discourse structure. Figure 2 provides an illustrative example, highlighting how structural flow modeling captures cross-turn dependencies often overlooked in previous work. (b) Intent inference: By analyzing structural patterns, this taxonomy facilitates the extraction of implicit user intent, offering a deeper understanding of how instructions evolve over multiple turns. (c) Controlled generation: The taxonomy provides configurable structural parameters that guide task-specific dialogue simulation, allowing for the tailored generation of multi-turn conversations with predefined structural patterns. This not only enhances dataset diversity but also supports the development of more robust instruction-following models adaptable to varied real-world applications.
@@
  • Structural Flow Taxonomy: We propose a six-category structural taxonomy for multi-turn instruction-following evaluation, offering an interpretable framework for analyzing dialogue structural flow.
  • StructFlowBench: We introduce StructFlowBench, a structurally annotated multi-turn benchmark that leverages a structure-driven generation paradigm to better simulate complex dialogue scenarios.
@@ 2.1 Benchmarks for Multi-Turn Dialogues
The evolution of dialogue evaluation paradigms has progressed from single-turn assessments to sophisticated multi-turn interaction analysis Wang et al. (2023); Sun et al. (2024); Duan et al. (2024). Among these, MT-Bench Zheng et al. (2023b) pioneered this transition by providing methodologies specifically designed to assess a model’s ability to handle multi-turn interactions. Building upon this, MT-Bench-101 Bai et al. (2024) introduces a more granular evaluation framework to assess fine-grained capabilities. Multi-IF He et al. (2024b) expands single-turn dialogues into multi-turn interactions by following simple, predefined linear paths. However, most existing work on multi-turn dialogue evaluation does not prioritize instruction following assessment and overlooks the role of structural information. MT-Eval Kwan et al. (2024) explores four types of multi-turn dialogue structures—recollection, expansion, refinement, and follow-up—which partially inspire our structural framework. However, MT-Eval does not establish a systematic structural framework and lacks integration of various structural aspects for a comprehensive evaluation.
@@ 2.2 Benchmarks for Instruction Following
Recent instruction following evaluation predominantly employs constraint-based frameworks Jiang et al. (2024); Zhang et al. (2024); He et al. (2024a); Zhou et al. (2023). Constraints are commonly used in these works to guide and evaluate LLM outputs, serving as criteria for assessing instruction-following behavior. In our work, we further extend this usage by employing constraints not only for evaluation but also as guidance for model generation. InfoBench Qin et al. (2024) introduces the Decomposed Requirements Following Ratio (DRFR) metric, which provides a more granular scoring system by breaking down the evaluation of complex instructions into assessments of their individual simple constraints. Furthermore, ComplexBench Wen et al. (2024) explores instruction-following capabilities in single-turn complex dialogues through empirical studies of constraint composition. However, prior work on instruction-following evaluation has primarily focused on single-turn interactions, which are less representative of real-world multi-turn usage. While some studies have attempted to split complex single-turn instructions into multi-turn dialogues, these approaches do not fully capture the intentionality and goal-oriented nature of users in real-world contexts.
@@ 3 StructFlowBench
This section introduces the structural flow framework and constraint categories, details the data construction pipeline and benchmark statistics, and outlines the evaluation protocol.
@@ 3.1 Structural Flow Taxonomy
By analyzing existing LLM and real human multi-turn dialogue datasets (such as WILDCHAT Zhao et al. (2024) and LMSYS-Chat-1M dataset Zheng et al. (2023a)), we identified and categorized six structural patterns of multi-turn dialogues to enhance the understanding and analysis of conversational structural flow. Descriptions of these six structures are illustrated in Figure 1.
Follow-up: An adjacent-turn structure where the user’s next prompt builds on the previous turn, incorporating details from either the user’s previous prompt or the AI’s previous response. This is the most common structure in multi-turn dialogues, typically reflecting the user’s intent to explore the topic more deeply.
Refinement: An adjacent-turn structure in which the user modifies or clarifies their immediate previous prompt to improve the AI’s response. This structure usually reflects dissatisfaction with the prior response and prompts the user to revise the prompt to better convey their concerns.
Recall: A long-range structure where the user refers back to content from two or more turns ago, either to provide context for the current prompt (long-range follow-up) or to seek clarification (long-range refinement).
@@
These six dialogue structures form the Structural Flow Taxonomy, which we use to analyze multi-turn dialogues and construct corresponding structural flows.
@@ 3.2 Constraint Categories
We categorize our constraints into intra-turn constraints and multi-turn structural constraints. The definitions and examples of all constraints are provided in Appendix Table 4, and their distribution is presented in Appendix Table 5.
For intra-turn constraints, we synthesize and refine constraint classification systems from several works in this field (e.g., IF-Eval Zhou et al. (2023), CFBench Zhang et al. (2024), FollowBench Jiang et al. (2024)). Based on this synthesis, we categorize constraints into eight types: Inverse Constraint, Style Constraint, Situation Constraint, Keyword/Element Constraint, Basic Format Constraint, Quantity Format Constraint, Template Format Constraint and Content Constraint.
For multi-turn structural constraints, we define five types of structural constraints, excluding the “unrelatedness” structure. These constraints are specifically designed to maintain logical coherence and continuity across multiple turns in a dialogue. They ensure that the structural relationships between turns are consistent and contextually relevant, enabling a smooth flow of conversation. The five types of constraints are aimed at handling key aspects such as follow-ups, refinements, recalls, expansions, and summaries, ensuring that each turn in the dialogue properly connects to the previous ones while adhering to the intended conversational structure.
@@ 3.3 Data Construction Pipeline
The construction pipeline of StructFlowBench, as shown in Figure 3, comprises three main components: parameter setting, two-step dialogue generation, and constraint extraction and addition. All prompt templates used in the data construction process are included in Appendix C, and a sample data instance is provided in Appendix Table 8.
@@
Before dialogue generation, we select parameters such as topic, task, user characteristics, and structural flow template, ensuring comprehensive coverage of the evaluation scope for multi-turn dialogue generation. For task types, we refer to the taxonomy of ComplexBench Wen et al. (2024), adapting it to our evaluation framework and selecting eight task types. For topics, we draw from the MT-Bench-101 Bai et al. (2024) framework, making necessary adjustments to suit our context, and ultimately select 22 topics. For user characteristics, we consider the significant differences in questioning styles and language between experts and non-experts. Regarding the structural flow templates, we manually designed 14 templates from real dialogue data (WildChat), covering most pairwise combinations of six structural relations.
@@
Table 1: Comparisons of StructFlowbench with existing benchmarks in terms of constraint types, multi-turn coverage, and structural information. $\triangle$ represents partially satisfied.
Two-Step Dialogue Generation
We employ a two-step process to generate a dialogue for a parameter setting. The first step uses the structural flow template to generate an intermediate dialogue plan (i.e., summarized prompts) via GPT-4o. The detailed prompt template is provided in Appendix Figure 7. Locally deployed mini-models perform initial screening and manual inspection of error data to ensure the dialogue plan aligns with the structural flow. In the second step, each intermediate dialogue plan is used to generate a complete dialogue, including user prompts and LLM responses via GPT-4o. The detailed prompt template is provided in Appendix Figure 8. This approach ensures high-quality generation of both dialogue content and structure while minimizing manual effort.
Constraint Extraction and Addition
For the complete multi-turn dialogue data, we extract intra-turn constraints using the GPT-4o, followed by manual validation to ensure accuracy. Further details are provided in Appendix Figure 9. Based on the structural flow information, we then assign the corresponding multi-turn structural constraints to each dialogue turn.
3.4 Benchmark Dataset Statistics
@@ 3.5 Evaluation
Deepseek-v3 0.99 0.8 0.92 1.0 1.0 0.97 0.93 0.96 0.98
Gemini-1.5-Pro 0.97 0.78 0.91 1.0 0.94 0.96 0.91 0.95 0.96
GPT-4o 0.98 0.78 0.88 0.97 0.91 0.96 0.9 0.95 0.96
Claude-3.5-Sonnet 0.98 0.8 0.88 1.0 0.91 0.95 0.89 0.94 0.95
GLM-4-9B-Chat 0.95 0.75 0.84 0.97 0.94 0.95 0.87 0.93 0.95
Qwen2.5-14B-Instruct 0.97 0.73 0.87 0.97 0.97 0.93 0.84 0.92 0.93
Qwen2.5-7B-Instruct 0.95 0.76 0.9 0.94 0.97 0.93 0.84 0.92 0.93
Deepseek-R1-Distill-Qwen-7B 0.91 0.62 0.85 0.86 0.78 0.81 0.7 0.8 0.82
@@
Llama-3.1-Instruct-8B 0.96 0.71 0.84 0.79 0.94 0.84 0.69 0.83 0.85
Phi-3.5-mini-instruct 0.94 0.68 0.87 0.94 0.94 0.88 0.74 0.87 0.87
@@
Mistral-7B-Instruct-v0.3 0.97 0.59 0.87 0.71 0.97 0.76 0.57 0.76 0.77
Table 2: StructFlowBench rated by GPT-4o. The left side of the figure displays the performance of various models on the five basic structural constraints, with accuracy used as the evaluation metric, while the right side presents their performance on the four key metrics.
@@ 4.1 Experimental Setup
We evaluate 13 popular LLMs on StructFlowBench, including 3 closed-source models (GPT-4o Hurst et al. (2024), Claude-3.5-Sonnet Anthropic (2024) and Gemini-1.5-Pro Team et al. (2024) ) and 10 open-source models: Llama-3.1-Instruct-8B Dubey et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct Yang et al. (2024), Yi-6B-Chat Young et al. (2024), Phi-3.5-mini-instruct Abdin et al. (2024), GLM-4-9B-Chat GLM et al. (2024), Deepseek-R1-Distill-Llama-8B, Deepseek-R1-Distill-Qwen-7B Guo et al. (2025) and DeepSeek-v3 Liu et al. (2024). More details on these evaluated models can be found in Appendix Table 9.